You are given the following information: The yield of a one year zero coupon bond is 2.0% while the yields on a 2 and 3 year fixed coupon bonds are 2.5% and 3.5% respectively. What is the 2-year Zero Rate?

Answer this :What is the expected one-year interest rate two years from today?
Answer #1
Formulae :
Forward rate=(1+rb​)tb/​(1+ra​)ta​​−1
Where:
ra​=The spot rate for the bond of term ( 3 Year) ta​ periods=3.5%
rb​=The spot rate for the bond with a ( 2 Year) shorter term of tb​ periods​=2.5%
= (1+0.035)3/(1+0.025)2-1
= 1.108718 -1
1.050625
= 5.53% i.e. one year interest rate two years from today.